OpenStudy (anonymous):

ASAP PLEASE SOMEBODY HELP ME NOW!!!!!!!!! Marci finished the one hundred yard dash one and four-fifths seconds faster than Eva. If Eva finished in fourteen and seven-tenths seconds, how long did it take Marci to finish?

OpenStudy (tkhunny):

Eva: 14.7 Marci: 14.7 - 1.8 There are two challenges, here. One is recognizing that Marci's time can be calculated by subtraction from Eva's time. The other is recognizing 4/5 = 8/10 = 0.8

OpenStudy (anonymous):

@tkhunny what are you saying? are you saying that the answer because im afraid not

OpenStudy (anonymous):

@Juliette32801 So Eva had a time of 14 9/10 and Marci finished 1 4/5 seconds later. First you need to calculate Marci's time so that its ten or tenths (i dunno). Then it becomes 1 8/10. Add 9/10 + 8/10 + 17/10. With is improper and then it becomes 1 7/10. So Marci finished 13 2/10 seconds. Which is equivalent to 13 1/5.

OpenStudy (anonymous):

*= 17/10

OpenStudy (anonymous):

@SchoolisCool are you saying 12 9/10 is the answer

OpenStudy (anonymous):

@Juliette32801 Is that correct?

OpenStudy (tkhunny):

I am NEVER saying the answer. Read the Code of Conduct for my reasons why. I am to provide hints and instruction. You tell me the answers that you determine.

OpenStudy (anonymous):

@SchoolisCool no but its close to one of my answer choices which is 16 1/2

OpenStudy (anonymous):

Well you didn't exactly give a good description @tkhunny

OpenStudy (anonymous):

@tkhunny I know you not saying the answer but you explain with the answer and by the way.. Code Of Conduct BE NICE!

OpenStudy (tkhunny):

16 1/2 is no good. That would mean that Marci was slower!

OpenStudy (anonymous):

@SchoolisCool yep

OpenStudy (anonymous):

No I'm saying is my answer. 13 2/10 or 13 1/5

OpenStudy (anonymous):

Im not sure

OpenStudy (anonymous):

@SchoolisCool close to that in my answer choices is 12 9/10

OpenStudy (tkhunny):

You cannot show where I am not being nice. Eva: 14.7 Marci: 14.7 - 1.8 Do the subtraction.

OpenStudy (anonymous):

you put NEVER in cap meaning your yelling (:

OpenStudy (tkhunny):

A common misunderstanding. I am only emphasizing. I NEVER yell. :-)

OpenStudy (anonymous):

@tkhunny mm sure.

OpenStudy (tkhunny):

It is important to have in your mind acceptable answers. If we know Eva is 14.7 and Marci is faster, we can immediately discard all possible answers greater than 14.7.

OpenStudy (anonymous):

@tkhunny I know the answer its 12 9/10

OpenStudy (tkhunny):

You know it because it was given to you or because you calculated it?

OpenStudy (anonymous):

why does it matter? @tkhunny

OpenStudy (anonymous):

@tkhunny neither I subtracted like you told me remember the code of conduct (:

OpenStudy (tkhunny):

If you subtracted like I told you, then you calculated it. Good work.